FLOSS01: Logbook Entries

FLOSS01: Kryptons Messages: 14 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
176 Mon 18-Mar-2002Kryptons bad at PAM sutessemmer
173 Wed 13-Mar-2002rad & kryptons cleanedall stnsoncley
169 Wed 13-Mar-2002krytpons cleanedtoweroncley
161 Fri 01-Mar-2002Replace Kr at Sagestn3delany
153 Fri 22-Feb-2002Kr and Rad cleanall stnsdelany
143 Tue 19-Feb-2002new kryptonstn2oncley
142 Tue 19-Feb-2002new kryptonstn3oncley
139 Mon 18-Feb-2002kryptons cleanedall stnsoncley
120 Fri 18-Jan-2002Added h2oV.barestn2maclean
115 Wed 16-Jan-2002Wrong cal used for krypton at sage sitestn3maclean
110 Fri 04-Jan-2002PAM Kr cleanedstn3delany
105 Fri 04-Jan-2002Kryptons Cleanedtowermilitzer
80 Sun 09-Dec-2001krypton cleanedstn2oncley
77 Sun 09-Dec-2001kryptons cleanedtoweroncley


176: Kryptons, Site , Mon 18-Mar-2002 08:38:58 MST, Kryptons bad at PAM sutes
The kryptons appear to be suffering from
low temperature problems at the PAM sites.
This has been noted before. The actual cause
has not been determined (As far as I know!)

173: Kryptons, Site all stns, Wed 13-Mar-2002 16:51:47 MST, rad & kryptons cleaned
Sites 2 first then 3.

169: Kryptons, Site tower, Wed 13-Mar-2002 15:51:25 MST, krytpons cleaned
Just cleaned kryptons at 10m and 2m on the tower.

161: Kryptons, Site stn3, Fri 01-Mar-2002 09:37:19 MST, Replace Kr at Sage
Inspected Sage site. All vent + asp fans OK
Kr #1133 was oriented with the receiver pointing to the S.
Kr 1133 was removed.
Kr 1395 was installed with receiver pointing to the N (wire S of attachment)
Had trouble sucuring the device but got screwws started and then used panduits.

153: Kryptons, Site all stns, Fri 22-Feb-2002 11:25:17 MST, Kr and Rad clean
At 10:50 visited Sage and cleaned Kr and all radiometers
At 11:00 visited Bare and cleaned Kr and netrad
At 11:10 visited tower and cleaned Kr.2m and all radiometers

Note in no case was there any indication of dirt on the domes.
143: Kryptons, Site stn2, Tue 19-Feb-2002 14:27:43 MST, new krypton
at bare, just swapped krypton 1101 for 1392, again to attempt to solve the
cold temperature problem.

142: Kryptons, Site stn3, Tue 19-Feb-2002 14:25:28 MST, new krypton
at sage, we just swapped 1133 for 1395 in hopes of solving the cold temperature
dropout problem

I also checked the fence battery which is now 14.5V connected through the 
solar panel, so I think all is well.

I also note the serial numbers:
csat 538
trh 703

139: Kryptons, Site all stns, Mon 18-Feb-2002 15:47:14 MST, kryptons cleaned
All kryptons have now been cleaned, starting with the top of the main tower at
about 13:00 and progressing down, the sage ~1500 and bare ~1530.

120: Kryptons, Site stn2, Fri 18-Jan-2002 08:28:56 MST, Added h2oV.bare
Last night we changed the setup at stn2 (bare site) to sample the
krypton voltage through the Eve A2D, in addition to the serializer.

The new variable is h2oV.bare - it is being covar'd into the netcdf files.


Last night, at least, the voltage reported by the serializer agrees
with h2oV - they were both around 2 V.


Note that the serializer voltage as reported with the Eve "KRYPTON"
command on the console appears to be wrong.  

The q_kh2o_serializer cal function converts serializer counts to volts with

  volts = counts * 5 / 32767

On a big-endian system (eve, and suns) you must swap the bytes
of the counts before the conversion.  Eve is either not swapping
bytes or is using some other counts->volts conversion.

**** Note this this problem only effects the display on the eve
console of krypton voltage from the serializer - it does not
effect the values computed by the base system. ****


115: Kryptons, Site stn3, Wed 16-Jan-2002 19:36:40 MST, Wrong cal used for krypton at sage site
Steve Oncley found the problem with the krypton at the sage site:

  prep.config is using cal_files/krypton1392 for both s2 and s3.
  s3 should have krypton1395.  s3 has been going below 0 since 
  its voltage has been about 3.8 V and Vo for 1392 was 3.6V. 
  Vo for 1395 is 4.9V.


This has now been corrected, and the netcdf data for the sage h2o
variable are being re-computed from the present back to Dec 17,
when it was installed.




110: Kryptons, Site stn3, Fri 04-Jan-2002 16:43:13 MST, PAM Kr cleaned
Cleaned Kr at the two PAM stations at 16-30

105: Kryptons, Site tower, Fri 04-Jan-2002 15:56:25 MST, Kryptons Cleaned
01/04/2002

13:35 2m
13:50 10m
14:00 20m

80: Kryptons, Site stn2, Sun 09-Dec-2001 17:03:17 MST, krypton cleaned
I finally got to the bare site to clean its krypton about 1600.
77: Kryptons, Site tower, Sun 09-Dec-2001 11:12:03 MST, kryptons cleaned
Larry and I just cleaned all the kryptons, starting at 20m over the last hour.
Of course, we had to pull up the booms at 20 & 10, so sonic data will be bad.